Calculate the mass of ethyl alcohol which must be dissolved in 750 grams of water to make 0.5m solution.
1
Expert's answer
2020-09-02T10:09:25-0400

The molar concentration "c" of ethyl alcohol solution is the ratio of the number of the moles of ethyl alcohol "n" and the volume of the solution "V".

"c =\frac{n}{V}" .

The density of 0.5 M solution of ethanol at 20°C is 0.9939 g/mL. The molar mass "M" of ethanol is 46.07 g/mol. Therefore:

"c = \frac{m_{eth}/M}{(m_w+m_{eth})/d}"

"0.5 = \frac{m_{eth}/46.07}{(750+m_{eth})/0.9939/1000}"

"m_{eth} = 17.8" g.

Answer: 17.8 g of ethyl alcohol must be dissolved in 750 grams of water to make 0.5M solution at 20°C.